OpenStudy (sette):

Algebra 1 Honors: Subtracting Rational Expressions what is the difference of (x/x-5)-[(3x-6)/(x-5)^2] ?

OpenStudy (anonymous):

common denominator will be \[(x-5)^2\]

OpenStudy (anonymous):

\[\frac{x}{x-5}-\frac{3x-6}{(x-5)^2}\] \[\frac{x(x-5)-(3x-6)}{(x-5)^2}\] then algebra to simplify the numerator

OpenStudy (sette):

OKay Thank You